### v1

# Gravitational Bound State Perturbations Inside Black Holes and Isospectrality

Abstract
We study the bound state solutions for the polar perturbations in the interior of the Schwarzschild black hole. It is shown that for a given value of the spherical harmonic index $\ell$, there are a total of $\ell-1$ bound states for polar perturbations. We show both analytically and numerically that the spectrum of $\ell-2$ of these perturbations agrees identically with the spectrum of axial perturbations. Correspondingly, the isospectrality between the bound states of the axial and polar perturbations in the interior of the black hole does hold. This is significant since the effective potential is singular at the center of the black hole, and the existence of the isospectrality is far from obvious. Furthermore, we demonstrate that the additional mode found in the spectrum of polar perturbations is the algebraically special mode, which is the ground state of polar perturbations as well.
Introduction: The unperturbed black hole (BH) solutions of the Einstein field equation are mathematically simple and yet elegant Chandrasekhar (1985). However, the observable physical effects appear at the level of perturbations, such as in the detection of GWs emission during the merging of BHs and the ringdown phase as observed in LIGO/Virgo/KAGRA collaborations Abbott et al. (2016, 2017, 2020); LIG (2025). A key role in BH physics is played by the BH event horizon which separates the interior region of the BH from its exterior region. In particular, as the interior region is causally disconnected from the exterior region, it is assumed that the dynamical events happening in the interior region are inaccessible to the exterior observers and hence irrelevant for gravitational and cosmological observations. While this notion is practically viable, nonetheless, the interior of the BH is part of the whole manifold, and a consistent treatment of BH perturbations requires that the perturbations in the interior of the BH to be studied as well.
The question of bound state perturbations in the interior of the Schwarzschild BH was originally studied in Firouzjahi (2018) and was revisited more recently in Steinhauer et al. (2025); Firouzjahi (2025). For other works concerning the interior of BH see also Fiziev (2006); Firouzjahi (2024). The bound states are interpreted as the solutions with an imaginary spectrum in which the perturbations are regular at the center of the BH, while their time-dependent profile falls off exponentially on the event horizon. It is important to note that these perturbations are independent of perturbations in the exterior region and are defined intrinsically in the interior of BHs. It is demonstrated in Firouzjahi (2025) that scalar, vector, and axial tensor perturbations all admit bound state solutions. Specifically, it is shown in Firouzjahi (2025) that for a given value of the angular index $\ell$ and the spin $s$ of the fields, there are a total of $\ell-s$ bound state solutions. Finally, it is shown in Steinhauer et al. (2025) that the bound state solutions inside the BHs have the curious property that their total wave function has a non-zero amplitude on the event horizon. Correspondingly, these bound states are named in Steinhauer et al. (2025) as overdamped quasibound states (OQBSs).
The quasi-normal modes (QNMs) perturbations in the exterior of BHs are vastly studied, for some reviews and an incomplete list of theoretical studies see Nollert (1999); Kokkotas and Schmidt (1999); Berti et al. (2009); Konoplya and Zhidenko (2011); Berti et al. (2025); Iyer and Will (1987); Motl and Neitzke (2003); Andersson and Howls (2004); Cardoso et al. (2003); Medved et al. (2004); Berti et al. (2006); Maggiore (2008); Konoplya et al. (2019); Hod (2025). The scalar, vector, and the axial tensor perturbations are governed by the Regge-Wheeler (RW) equation, while the corresponding equation for the polar tensor perturbations is the Zerilli equation Regge and Wheeler (1957); Zerilli (1970); Chandrasekhar (1985). As the Zerilli potential is more complicated than the RW potential, the bound state perturbations of polar perturbations were not studied in Steinhauer et al. (2025) and Firouzjahi (2025). In the exterior region of BHs, there is a one-to-one transformation which links the polar perturbations to the axial perturbations Chandrasekhar (1985); Chandrasekhar and Detweiler (1975); Anderson and Price (1991). Since this transformation is regular in the exterior region, an immediate conclusion is that the axial and polar perturbations have the same spectra of QNMs perturbations.
In this work, we extend upon the analysis of Steinhauer et al. (2025); Firouzjahi (2025) and study the bound state perturbations for the polar perturbations in the interior of the Schwarzschild BH. It is important to note that the bound state solutions considered here and in Firouzjahi (2018); Steinhauer et al. (2025); Firouzjahi (2025) are different from the bound state solutions of the inverted potential proposed by Mashhoon and collaborators Mashhoon (1982); Blome and Mashhoon (1981); Ferrari and Mashhoon (1984a, b); Liu and Mashhoon (1996). More specifically, the bound states in Mashhoon (1982); Blome and Mashhoon (1981); Ferrari and Mashhoon (1984a, b); Liu and Mashhoon (1996) are obtained from the complex transformation of the RW equation to bring the potential in the form of Poschl-Teller or the Eckart potentials which can be solved analytically. This method was further employed recently in Völkel (2025) (see also Hatsuda (2020)) in which the inverted RW potential is considered directly and a numerical method is used to determine the eigenvalues of energies. In contrast, the bound states in our study are interpreted as the perturbations in the interior of the BH with imaginary frequencies such that their profiles are regular and confined to the interior of the BH which are inaccessible to an exterior observer. With these regularity boundary conditions, the profiles of the bound states are square integrable in the interior of the BH.
Gravitational Perturbations Inside the BH: The Schwarzschild metric in spherical coordinates is given as usual by,
[…]
in which $V_{\mathrm{eff}}$ takes either the RW or the Zerilli potentials denoted by $V^{-}$ and $V^{+}$ respectively. Similar to Steinhauer et al. (2025); Firouzjahi (2025), we look for bound state solution $\omega=i\,\omega_{I}$ with $\omega_{I}>0$.
As it is well-known, there are relations between $V^{\pm}$ and between the corresponding mode functions $Z^{\pm}$ which ensure the isospectrality of the QNM spectrum associated to the axial and polar perturbations in the exterior region Chandrasekhar (1985); Chandrasekhar and Detweiler (1975). More specifically,
[…]
and
$\Big(4\lambda(\lambda+1){\mp}6i\omega\Big)Z^{\pm}$ $=$ $\Big[4\lambda(\lambda+1)+\frac{18(r^{2}-r)}{r^{3}(2\lambda r+3)}\Big]Z^{\mp}$ (4)
[…]
in which $2\lambda\equiv(\ell-1)(\ell+2)$ and,
$\hskip 28.45274ptf\equiv\frac{r-1}{r^{2}(2\lambda r+3)}\,.$ (5)
Indeed, one can specifically check that starting with $Z^{\pm}$ as the solution of the master equation (2) and using the transformations (3) and (4), then $Z^{\mp}$ satisfies Eq. (2) as well.
The specific forms of $V^{\pm}$ are presented in Appendix A while their plots are shown in Fig. 1. It is seen that $V^{-}$ develops a negative global minimum before diverging like $V^{-}\sim 3/r^{4}$ near the center. On the other hand, $V^{+}$ is always negative and diverges like $V^{+}\sim-1/r^{4}$ at $r=0$. The fact that $V^{-}$ develops a negative global minimum indicates that the bound states with $\omega=i\,\omega_{I}$ can exist for axial perturbations Steinhauer et al. (2025); Firouzjahi (2025) . However, the existence of bound states for $V^{+}$ may look non-trivial. Indeed, we confirm that the bound states for polar perturbations exist as well.
Intuitively, the transformations (3) and (4) may look singular at $r=0$. More specifically, the regularity of the bound state solutions at $r=0$ requires that $Z^{-}\sim r^{3}$ while $Z^{+}\sim r$. Plugging these asymptotic forms in (3) and (4) one may conclude that Steinhauer et al. (2025) $Z^{-}_{\mathrm{reg}}\rightarrow Z^{+}_{\mathrm{reg}}$ while $Z^{+}_{\mathrm{reg}}\rightarrow Z^{-}_{\mathrm{sing}}$. If so, this implies that starting with the regular axial perturbations one can obtain the regular polar perturbation, but not the other way around. However, this is subtle and misleading. We will show that the transformations (3) and (4) are indeed regular and, except for one particular mode, starting with regular $Z^{\pm}$ one obtains the regular $Z^{\mp}$. The only exception is the “algebraically special mode” (ASM) which, for a given value of $\ell$, turns out to be the ground state of the polar perturbations.
Figure 1: The effective potentials $V^{-}(r_{*})$ and $V^{+}(r_{*})$ given in Eqs. (17) and (18) with $\ell=3$.
The ASM is the solution in which the left and right hand sides of Eq. (4) vanish simultaneously. With $\omega_{I}>0$, the left hand side of Eq. (4) allows only the lower sign associated to the polar perturbations with the spectrum Chandrasekhar (1985, 1984); Couch and Newman (1973); Wald (1973),
$\hskip 28.45274pt\omega_{\mathrm{sp}}=\frac{i}{6}(\ell-1)\ell(\ell+1)(\ell+2)=\frac{2i}{3}\lambda(1+\lambda)\,.$ (6)
On the other hand, solving the equation in the right hand side of (4), the profile of the polar perturbations associated to ASM is obtained to be,
$Z^{+}_{\mathrm{sp}}(r)=N_{+}(\lambda)\frac{r(1-r)^{-i\omega_{\mathrm{sp}}}}{3+2\lambda r}e^{-i\omega_{\mathrm{sp}}r}\,$, (7)
in which $N_{+}(\lambda)$ is a normalization constant. One can easily check that $Z^{+}_{\mathrm{sp}}(r)$ satisfies the Zerilli equation (2). Of course, the second order differential equation (2) has another solution for the polar perturbations. But, constructing the second solution via the Wronskian method, we have checked specifically that this second solution is singular at the horizon $r=1$ and is not allowed as a consistent bound state solution. On the other hand, since $Z^{+}_{\mathrm{sp}}(r)$ satisfies Eq. (4) identically at $\omega=\omega_{\mathrm{sp}}$, there is no equation left for $Z^{-}_{\mathrm{sp}}(r)=0$, meaning that the ASM is not shared by the axial perturbations. This is consistent with the results of Firouzjahi (2025) in which no bound state solutions for axial perturbations at the frequency $\omega=\omega_{\mathrm{sp}}$ was obtained.
In the next section we confirm that for all other bound states (i.e. $\omega\neq\omega_{\mathrm{sp}}$) the transformations (3) and (4) are regular. An immediate conclusion is that the isospectrality between the axial and polar bound state perturbations does hold. We confirm this interesting result by direct numerical analysis as well.
[…]
Figure 2: The profile of the normalized algebraically special mode $Z^{+}_{\mathrm{sp}}(r)$ given in Eq. (7). Curves from bottom to top correspond to $\ell=2,3,4,5$ respectively.
Isospectrality from SUSY Analysis: Here, using the supersymmetry (SUSY) method, we analytically confirm the existence of isospectrality between the polar and axial perturbations in the interior of BH and highlight the special role which is played by the ASM mode. We comment that this method is formally parallel to corresponding analysis for the QNMs of the exterior region, see for example Cardoso and Lemos (2001); Konoplya and Zhidenko (2011); Berti et al. (2009). However, there are important differences between the two setups which require independent investigations. First, we look for the bound state solutions which have different boundary conditions than QNMs in the exterior region. Second, the potentials $V^{\pm}$ are singular at $r=0$ so the regularity of transformations (3) and (4) is not manifest a priori Steinhauer et al. (2025). Here we present the main results while further technical details are provided in Appendix B. For a review of the SUSY method see Gangopadhyaya et al. (2017); Cooper et al. (1995).
Consider the superpotential,
$\mathcal{W}_{\ell}(r)\equiv\frac{3\,(1-r)}{r^{2}(3+2\lambda r)}-\frac{2\lambda(\lambda+1)}{3}=-3f+i\omega_{\mathrm{sp}}\,$, (8)
in the domain $0\leq r\leq 1$. Equipped with this superpotential, SUSY ladder operators are constructed as follows,
$\hat{\mathcal{D}}^{\pm}_{\ell}(r_{*})\equiv\pm\dfrac{d}{dr_{*}}+\mathcal{W}_{\ell}(r_{*})\,.$ (9)
Using the above ladder operators, we construct the partner Schrodinger Hamiltonians,
[…]
with the partner potentials given as,
[…]
The Schrodinger-like equations associated to this system are,
$\hskip 28.45274pt\mathcal{H}_{\ell}^{\pm}(r_{*})\phi^{\pm}_{\ell,n}=\mathcal{E}^{\pm}_{\ell,n}\phi^{\pm}_{\ell,n},\qquad n\geq 0,$ (12)
[…]
With $\mathcal{W}_{\ell}$ given in Eq. (8), one can easily show that the partner potentials $U^{\pm}_{\ell}$ are related to the original axial and polar potentials via $U^{\pm}_{\ell}=V^{\pm}-\omega_{\mathrm{sp}}^{2}$. Correspondingly, the Schrodinger-like equation (12) is equivalent to the original Eq. (2) with $\big(\mathcal{E}^{\pm}_{\ell,n}\big)^{2}=\big(\omega^{\pm}_{\ell,n}\big)^{2}-\omega_{\mathrm{sp}}^{2}$. Note that $\omega^{\pm}_{\ell,n}$ represents the spectrum of polar and axial perturbations at the level $n$. Furthermore, the values of the superpotential at the end points of its domain have opposite signs, $\mathcal{W}_{\ell}(0)=+\infty$, and $\mathcal{W}_{\ell}(1)<0$, so the SUSY is unbroken Gangopadhyaya et al. (2017) .
As mentioned before, the equation $\hat{\mathcal{D}}^{\pm}_{\ell}\,\phi^{\pm}_{\ell,0}=0$ specifies the ground states of these potentials with $\mathcal{E}^{\pm}_{\ell,0}=0$. Correspondingly, the ground states associated to each Hamiltonian are obtained to be,
$\phi^{+}_{\ell,0}(r)\sim\dfrac{1}{\phi^{-}_{\ell,0}(r)}\sim\dfrac{r\,(r-1)^{-i\omega_{\mathrm{sp}}}}{(3+2\lambda r)}\,e^{-i\omega_{\mathrm{sp}}\,r}\,.$ (13)
The ground state $\phi^{+}_{\ell,0}(r)$ is regular and normalizable in the domain of the superpotential while the state $\phi^{-}_{\ell,0}$, is singular at $r=0,1$ and non-normalizable; this confirm that the SUSY is unbroken.
As $U^{\pm}_{\ell}$ are partner potentials and SUSY is not broken, the isospectrality between their spectra does hold Gangopadhyaya et al. (2017). However, the ground state of $U^{+}_{\ell}$ can not be obtained from any of the eigenfunction of $U^{-}_{\ell}$ by operation of $\hat{\mathcal{D}}_{\ell}^{-}$, and it does not have any correspondence in the spectrum of $U^{-}_{\ell}$.
Since $\mathcal{H}_{\ell}^{\pm}$ are SUSY partners, their eigenfunctions and eigenvalues for $n=1,2,\dots$, are related to each other by,
$\phi^{\mp}_{\ell,n}(r_{*})=\dfrac{1}{\sqrt{\mathcal{E}^{\pm}_{\ell\,n\pm 1}}}\hat{\mathcal{D}}^{\pm}_{\ell}(r_{*})\,\phi^{\pm}_{\ell,n\pm 1}(r_{*}),$ (14)
$\mathcal{E}^{-}_{\ell\,n-1}=\mathcal{E}^{+}_{\ell\,n}$. (15)
Here the level number $n$ for $\phi^{-}_{\ell,n}$ is redefined such that the lowest normalizable sate of $\mathcal{H}_{\ell}^{-}$ is $\phi^{-}_{\ell,0}$.
As mentioned, the Schrodinger-like equation (12) is equivalent to the original equation (2). Consequently, the eigenfunctions of polar perturbations are given by $Z^{+}_{\ell,n}=\phi^{+}_{\ell,n}$. More specifically, the ground state is given by $Z^{+}_{\ell,n=0}(r)=\phi^{+}_{\ell,0}(r)$ with $\phi^{+}_{\ell,0}(r)$ as given in Eq. (13) with the spectrum $\omega^{+}_{\ell,n=0}=\omega_{\mathrm{sp}}$. Furthermore, the eigenfunctions of axial perturbations are given by $Z^{-}_{\ell,n}=\phi^{-}_{\ell,n}$. Finally, from Eqs. (14) and (15) and the fact that there are only $\ell-2$ axial bound states, we obtain,
$Z^{\mp}_{\ell,n}(r_{*})=\dfrac{1}{\sqrt{(\omega^{\pm}_{\ell\,n\pm 1})^{2}-\omega^{2}_{sp(\ell)}}}\hat{\mathcal{D}}^{\pm}_{\ell}(r_{*})Z^{\pm}_{\ell,n\pm 1}(r_{*})\,$, (16)
with $\omega^{+}_{\ell,n}=\omega^{-}_{\ell,n-1}$ for $n=1,2,\dots,\ell-2$.
The above analysis confirms that the isospectrality between polar and axial gravitational perturbations is preserved in the Schwarzschild interior. This isospectrality arises because $U^{\pm}_{\ell}$ are partner potentials in SUSY sense and the singularity of the potentials at $r=0$ does not destroy this property. In addition, the polar perturbations contain an extra mode that does not have any correspondence in axial modes. This is nothing but the ASM $Z^{+}_{\ell,n=0}=Z^{+}_{\ell,\mathrm{sp}}$ with the eigenfrequency $\omega_{\mathrm{sp}}$.
Numerical Results: Our numerical results also confirm the isospectrality and the existence of the ASMs. We present the detail numerical analysis in the Appendix A. Below is the summary of our main numerical results.
(a): For a given value of $\ell$, there are total $\ell-1$ bound state solutions for the polar perturbations. This should be compared with the case of axial perturbations, which admit $\ell-2$ solutions Firouzjahi (2025). (b): All $\ell-2$ spectra of the axial perturbations are identically shared with the polar perturbations. In this view, the isospectrality between the polar and axial perturbations does hold. We have confirmed this conclusion to 40 digits in our numerical analysis. (c): The additional mode in polar perturbations is indeed the ASM with the spectrum given in Eq. (6). This is the ground state solution for the polar perturbations in which, for a given value of $\ell$, $\omega^{2}$ has the least value (i.e., the most negative value). In particular, for $\ell=2$, $\omega_{\mathrm{sp}}$ is the only allowed spectrum. We have confirmed that the spectrum of ASM agrees with Eq. (6) to 40 digits. This is significant, noting that there are controversies on the exact numerical values of ASMs for QNMs Berti et al. (2009); Maassen van den Brink (2000); Leaver (1985); Nollert (1993); Liu and Mashhoon (1996). (d): The level number of the excited states with $n=1,2,$... is the same as the number of the nodes for the profile of $Z(r)$. For example, the first excited state has one node, the second excited state has two nodes, and so on. On the other hand, the most excited state with $n=\ell-2$ has $\ell-2$ nodes. (e): For large $\ell$ and for low-lying states (say $n=1$), the spectrum satisfies an analytic WKB formula as given in Firouzjahi (2025) in which $\omega_{I}\propto(\ell+\frac{1}{2})^{4}$. (f): For $\ell\gg 1$, the most excited state, which for a given value of $\ell$ has the lowest value of $\omega_{I}$, satisfies $\omega_{I}=1+c\,\ell^{-1}+{\cal O}(\ell^{-2})$ with $c\simeq 1$. An important conclusion from the above results is that independent of $\ell$ and $n$, all modes satisfy the lower bound $\omega_{I}>1$. Note that the conclusions (d), (e) and (f) listed above are identical to axial perturbations obtained in Firouzjahi (2025). Furthermore, conclusions (b) and (c) are confirmed analytically from our SUSY analysis as reviewed in previous section.
$n$ $\ell=2$ $\ell=3$ $\ell=4$ $\ell=5$ $\ell=6$ $\ell=7$
0 4.000000000 20.00000000 60.00000000 140.0000000 280.0000000 504.0000000
1 - 1.705649856 5.745972088 13.76767450 27.78537685 50.20490178
2 - - 1.394461407 3.968693503 8.389356734 15.42616838
3 - - - 1.274578721 3.315604823 6.487859859
4 - - - - 1.211013017 2.982338127
5 - - - - - 1.171550415
Table 1: $\omega_{I}$ for polar perturbations for different values of $\ell$. For each $\ell$, there are $\ell-1$ bound states in which the ASM is the ground state with the largest value of $\omega_{I}$ (most negative value of $\omega^{2}$) while the most excited state with $n=\ell-2$ has the lowest value of $\omega_{I}$. Here and in Fig. 3, $\omega_{I}$ are normalized in units of $(2GM)^{-1}$.
[…]
In Table 1, we have presented $\omega_{I}$ for some values of $\ell$. We have calculated the spectrum in 40 digits but in order to save space, here we only present the spectrum to 10 digits. One can check that, except for $\omega_{\mathrm{sp}}$ associated to the ground state $n=0$, all the remaining $\omega_{I}$ are the same as for axial perturbations presented in Firouzjahi (2025).
[…]
Finally, it is helpful to examine the similarities and differences of our bound states with QNMs in the exterior region. In both cases, the isospectrality is valid and the polar perturbations have the additional ASM excitations. However, a key difference is that for a given value of $\ell$ the number of bound states in the interior of the BH is finite but the number of QNMs is infinite. This is perhaps due to differences between the asymptotic shapes of the corresponding effective potentials in the interior and exterior regions. More specifically, $V^{-}_{\mathrm{eff}}$ can be approximated by Coulomb potential in far regions Liu and Mashhoon (1996). However, near the horizon in the interior $(r_{*}\rightarrow-\infty)$ it falls off exponentially like $e^{r_{*}}$ Firouzjahi (2018) (see Appendix A). Nonetheless, the question of why the number of bound states in the interior of BH is finite (for a given value of $\ell$) is important which deserves careful investigations. Another difference is that ASM in our analysis is the ground state ($n=0$), while in the spectrum of QNMs the ASM is far from the ground state. For example, for $\ell=2,3$, ASM corresponds to $n=9$ and $n=41$ in QNMs spectrum respectively Leaver (1985).
Conclusions: In this work, we have studied the bound states of the polar perturbations in the interior of the Schwarzschild BH and examined the isospectrality. The bound states are solutions with imaginary spectrum which are regular at the center of BH while their profiles fall off exponentially near the event horizon.
We have shown numerically that for a given value of $\ell$, there are $\ell-1$ bound states of polar perturbations in which the spectra of $\ell-2$ of these bound states are identical (within 40 digits in our numerical analysis) with those of axial perturbations. In this sense, the isospectrality between these two perturbations does hold. However, there is an additional mode in the spectrum of polar perturbations which turns out to be the ASM with the spectrum Eq. (6) and the mode function given in Eq. (7). The isospectrality between the polar and axial perturbations is a direct consequence of the map that links these two perturbations as we studied via SUSY method. The fact that the potentials are singular at $r=0$ does not destroy this correspondence. Unlike the QNMs spectrum, the number of bound states (for a given value of $\ell)$ is finite and the ASM is the ground state. As in Firouzjahi (2025), the bound states satisfy the lower bound $\omega>1$ which is saturated only asymptotically for $\ell\rightarrow\infty$. In Tables 1 we have presented the numerical results for $\omega_{I}$ for some values of $\ell$ while the profiles of the mode functions are presented in Figs. 2 and 3. From Fig. 3 we observe that the number of nodes is equal to the level number $n$.
Acknowledgment: We thank Bahram Mashhoon and Antonio Riotto for helpful discussions and correspondences. The work of H. F. is partially supported by INSF of Iran under the grant number 4046375.
Supplemental Material
Appendix A Numerical Analysis
Here we present the numerical analysis in some details.
The goal is to solve Eq. (2) numerically for polar perturbations, in which the axial and polar potentials are given by respectively,
$V^{-}(r_{*})=\big(1-\frac{1}{r}\big)\Big(\frac{\ell(\ell+1)}{r^{2}}-\frac{3}{r^{3}}\Big)\,$, (17)
and,
$V^{+}(r_{*})=\frac{r-1}{r^{4}}\Big[\frac{8\lambda^{2}(\lambda+1)r^{3}+12\lambda^{2}r^{2}+18\lambda r+9}{(3+2\lambda r)^{2}}\Big]$ (18)
[…]
One can check that near $r=0$, $V^{-}\simeq 3/r^{4}\simeq 3/4r_{*}^{2}$ while $V^{+}\simeq-1/r^{4}\simeq-1/4r_{*}^{2}$. On the other hand, noting that near the horizon $r-1\simeq-e^{r_{*}}$, one can show that,
$\hskip 28.45274ptV^{-}(r_{*}\rightarrow-\infty)\simeq(1-2\lambda)e^{r_{*}}\,$, (19)
and
$\hskip 28.45274ptV^{+}(r_{*}\rightarrow-\infty)\simeq-\frac{4\lambda^{2}+4\lambda+3}{3+2\lambda}e^{r_{*}}\,.$ (20)
The plots of $V^{\pm}(r_{*})$ given in Eqs. (17) and (18) for $\ell=3$ are presented in Fig. 1.
Employing the regularity of the wave function at $r=0$ and the asymptotic form of the bound state at $r=1$, and similar to Eq. (7) for ASM, we consider the following ansatz for the general polar perturbations,
$Z^{+}(r)\equiv\frac{r(r-1)^{-i\omega}e^{-i\omega r}}{\left(\lambda r+\frac{3}{2}\right)^{2}}g(r)\,$, (21)
in which the auxiliary function $g(r)$ is regular everywhere inside the BH.
Rewriting the corresponding Zerilli equation (2) in $r$ coordinates and plugging back the ansatz Eq. (21) in Eq. (2), we obtain the following equation for the auxiliary function $g(r)$,
[…]
$-2\bigg[4\lambda^{3}r^{2}+4\lambda^{2}r\left(3-i\omega r^{2}\right)+3\lambda(4r-1)+9i\omega r\bigg]g=0\,.$ (22)
As mentioned before, the function $g(r)$ is regular in the interval $0<r<1$ so we can use the following Frobenius series expansion,
$g(r)=\sum_{m=0}^{\infty}c_{m}r^{m}\,.$ (23)
Substituting the above series expansion into Eq. (22) leads to the following five-term recursion relation $c_{m},c_{m-1},c_{m-2},c_{m-3},c_{m-4}$,
$c_{m}=\frac{1}{3m^{2}}\Big[(3-4\lambda)m^{2}+(16\lambda-3)m-10\lambda\Big]c_{m-1}$
$-\frac{2}{9m^{2}}\Big[2\lambda^{2}\left(m^{2}-8m+18\right)-6\lambda\left(m^{2}-5m+4\right)$
$+9i\omega(m-1)\Big]c_{m-2}+\frac{4\lambda}{9m^{2}}\Big[\lambda\left(m^{2}-9m+18\right)$
$-6i\omega(m-3)-2\lambda^{2}\Big]c_{m-3}-\frac{8i\omega\lambda^{2}(m-5)}{9m^{2}}c_{m-4}\,$, (24)
with the following additional relations between the coefficients $c_{1},c_{2},c_{3}$ and $c_{0}$,
$c_{1}$ $=\frac{2}{3}\lambda c_{0}\,$, (25)
$c_{2}$ $=-\frac{1}{6}\left(2\lambda^{2}+2\lambda+3i\omega\right)c_{0}\,$, (26)
$c_{3}$ $=-\frac{1}{9}(\lambda+1)\Big(2\lambda(\lambda+1)+3i\omega\Big)c_{0}\,.$ (27)
Our five-term recursion relation here should be compared with corresponding three-term recursion relation for axial perturbations Leaver (1985); Nollert (1993); Firouzjahi (2025). This is because the Zerilli equation is algebraically more complicated than the RW equation.
For any value of $m\geq 4$, the recursion relation (24) and the conditions ̵(25)-(27) give a polynomial expression for $c_{m}$ in terms of $\omega_{I}$ with a normalization coefficient proportional to $c_{0}$. As supposed, the function $g(r)$ is regular at $r=0$ and in order for this function to be regular at $r=1$ too, it is necessary that the sum $\sum_{m=0}^{\infty}c_{m}$ to exist and to be finite. This criterion determines the spectrum of our bound state perturbations. Note that we are looking for the bound state solution with an imaginary spectrum $\omega=i\omega_{I}$ with $\omega_{I}>0$.
We use the numerical method prescribed in Firouzjahi (2018, 2025) to calculate the bounded-state spectrum. As argued in Firouzjahi (2018), in order for the sum $\sum_{m=0}^{\infty}c_{m}$ to exist and to be finite, it is sufficient that the coefficient $c_{m}$ becomes zero for a large value of $m$. If this condition holds for a large value of $m$, then the successive terms $c_{m+1}$, $c_{m+2}$, … are suppressed and we can expect that the sum $\sum_{m=0}^{\infty}c_{m}$ to converge quickly. The high accuracy of this method compared to the method of Leaver Leaver (1985) was confirmed in Firouzjahi (2018). Furthermore, our smooth numerical plots for the profile of $Z(r)$ as presented in Fig. 3 confirm that this approach is reliable. In addition, the fact that our numerical method predicts the existence of the ASM as a ground state solution, which is confirmed analytically as well, indicates that our numerical prescription is based on a solid ground.
Moreover, we have examined the validity of our numerical prescription described above by employing another independent numerical approach which is available online as well [^1] https://github.com/krezazadeh/bound_states_polar. This code works almost identically to Völkel’s method Völkel (2025), which is based on the computation of the Wronskian $W(r)=Z_{\rightarrow}^{+}\frac{dZ_{\leftarrow}^{+}}{dr}-\frac{dZ_{\rightarrow}^{+}}{dr}Z_{\leftarrow}^{+}$, where $Z_{\rightarrow}^{+}$ and $Z_{\leftarrow}^{+}$ are the solutions of the Zerilli differential equation (in $r$ coordinate) by integrating from the right and left endpoints of the interval $[0,1]$ respectively. To calculate these functions, we applied the eighth-order Runge-Kutta (RK8) method in our code to solve the differential equation (22) by using the series (23) as boundary conditions. Our code employs a random-walk minimization method to minimize $|W|$ evaluated within the integration interval. The code is written on the basis of the message passing interface (MPI) method, which allows parallel processing and speeds up the minimization process significantly. Starting from a guess value for $\omega$ that is within the radius of convergence of the desired root, we calculate the value of $|W|$. Then, we consider a value close to the previous frequency and calculate $|W|$ for it as well. If the result for the new frequency is less than the corresponding value for the previous frequency, we search around the new frequency; otherwise, we return to the previous one. We repeat this procedure until we reach the desired accuracy for the bound state frequency. We have checked that the results obtained from this independent method confirm our numerical results with very good accuracy.
Appendix B SUSY-Inspired Methods:
In this section, we briefly review SUSY approach for analyzing the isospectrality associated to a general one-dimensional superpotential. For further reviews on SUSY approach see Gangopadhyaya et al. (2017); Cooper et al. (1995).
Consider a real function $\mathcal{W}_{\ell}(x)$ which is called the superpotential. Using $\mathcal{W}_{\ell}(x)$, the conjugate ladder operators are defined on the physical Hilbert space as follows,
$\hat{\mathcal{D}}_{\ell}^{\pm}(x)=\pm\dfrac{d}{dx}+\mathcal{W}_{\ell}(x).$ (28)
It is straightforward to define the Hermitian operators,
$\mathcal{H}_{\ell}^{\pm}(x)=\hat{\mathcal{D}}_{\ell}^{\mp}(x)\hat{\mathcal{D}}_{\ell}^{\pm}(x)=-\dfrac{d^{2}}{dx^{2}}+U_{\ell}^{\pm}(x)\,.$ (29)
These operators have the familiar form of the Schrodinger Hamiltonian with the potentials given in term of the superpotential $\mathcal{W}_{\ell}(x)$ as follows,
$U_{\ell}^{\pm}(x)=\mathcal{W}_{\ell}^{2}(x)\mp\dfrac{d\mathcal{W}_{\ell}(x)}{dx}\,.$ (30)
In this view $\mathcal{H}_{\ell}^{\pm}(x)$ are called the partner Hamiltonians and the $U_{\ell}^{\pm}(x)$ are the partner potentials. These Hamiltonian possesses finite or infinite set of physical normalizable bound states which form the basis for the physical Hilbert spaces.
The physical normalizable and bounded eigenfunctions and eigenvalues of $\mathcal{H}_{\ell}^{\pm}(x)$ are denoted by $\phi^{\pm}_{\ell n}(x)$ and $\mathcal{E}^{\pm}_{\ell n}$ for $n=0,1,2,\dots$ (this sequence can be finite or infinite). It is straightforward to show that,
  • Due to the forms of $\hat{\mathcal{D}}_{\ell}^{\pm}(x)$ and their relations to $\mathcal{H}_{\ell}^{\pm}(x)$ then $\mathcal{E}^{\pm}_{\ell n}\geq 0$.
  • If $\hat{\mathcal{D}}_{\ell}^{\pm}(x)\phi^{\pm}_{\ell 0}(x)=0$ then $\mathcal{H}_{\ell}^{\pm}(x)\phi^{\pm}_{\ell 0}(x)=0$, which means that $\phi^{\pm}_{\ell 0}(x)$ is eigenvector of $\mathcal{H}_{\ell}^{\pm}(x)$ with eigenvalue zero $\mathcal{E}^{\pm}_{\ell 0}=0$ (ground state). Furthermore, the ground state eigenfunction can be written as $\phi^{\pm}_{\ell 0}(x)=N^{\pm}\,e^{\mp\int_{x_{0}}^{x}\mathcal{W}_{\ell}(x)dx}$, in which $N^{\pm}$ are the normalization constants.
It is not possible to simultaneously normalize the ground states $\phi^{\pm}_{\ell 0}(x)$ due to the relation $\phi^{+}_{\ell 0}(x)\sim\frac{1}{\phi^{-}_{\ell 0}(x)}$. This puts strong restrictions on the system. In the case in which $\phi^{\pm}_{\ell 0}(x)$ is normalizable and $\phi^{\mp}_{\ell 0}(x)$ is not, the ground energy of $\mathcal{H}_{\ell}^{\pm}(x)$ is zero and it is said that the SUSY is unbroken. In the case in which none of $\phi^{\pm}_{\ell 0}(x)$ are normalizable, neither of the two Hamiltonians have a zero-energy bound state, and it is said that the SUSY is spontaneously broken. It is possible to show that a superpotential that has different signs (the same sign) at the ends of the domain will generate a system with unbroken SUSY (spontaneously broken SUSY) Gangopadhyaya et al. (2017).
In the case of unbroken SUSY, we suppose the Hamiltonian that possess the normalizable ground state is $\mathcal{H}_{\ell}^{+}(x)$. As the two Hamiltonians $\mathcal{H}_{\ell}^{\pm}(x)$ originate from the same superpotential they share some common properties, like isospectrality. This is defined in Gangopadhyaya et al. (2017) as both Hamiltonians $\mathcal{H}_{\ell}^{\pm}(x)$ having the same set of eigenvalues, except possibly for the ground state of $\mathcal{H}_{\ell}^{+}(x)$ if this ground state exists.
By representing the eigenfunctions and the corresponding eigenvalues of $\mathcal{H}_{\ell}^{\pm}(x)$ by $\phi^{\pm}_{\ell n}(x)$ and $\mathcal{E}^{\pm}_{\ell n}$ for $n=1,2,\dots$ (note, $n\neq 0$), and using the definition of $\mathcal{H}_{\ell}^{\pm}(x)$, one can show that,
$\mathcal{H}_{\ell}^{\mp}(x)\left(\hat{\mathcal{D}}_{\ell}^{\pm}(x)\phi^{\pm}_{\ell n}(x)\right)=\mathcal{E}^{\pm}_{\ell n}\left(\hat{\mathcal{D}}_{\ell}^{\pm}(x)\phi^{\pm}_{\ell n}(x)\right)\,.$ (31)
This shows that $\hat{\mathcal{D}}_{\ell}^{\pm}(x)$ transfers the eigenfunction $\phi^{\pm}_{\ell n}(x)$ of $\mathcal{H}_{\ell}^{\pm}(x)$ with the eigenvalue $\mathcal{E}^{\pm}_{\ell n}$ to the eigenfunction of $\mathcal{H}_{\ell}^{\mp}(x)$ with the same eigenvalue $\mathcal{E}^{\pm}_{\ell n}$.
In the case that SUSY is not broken, it can be shown that Gangopadhyaya et al. (2017),
$\phi^{\mp}_{\ell n}(x)=\dfrac{1}{\sqrt{\mathcal{E}^{\pm}_{\ell n\pm 1}}}\hat{\mathcal{D}}_{\ell}^{\pm}(x)\phi^{\pm}_{\ell n\pm 1}(x),\quad\mathcal{E}^{-}_{\ell n-1}=\mathcal{E}^{+}_{\ell n}$, (32)
for $n=1,2,3\dots$. The number of $n$ for $\phi^{-}_{\ell n}(x)$ is redefined such that the lowest normalizable sate of $\mathcal{H}_{\ell}^{-}(x)$ to be $\phi^{-}_{\ell 0}(x)$. Furthermore, the ground state of $\mathcal{H}_{\ell}^{+}$ satisfies $\hat{\mathcal{D}}_{\ell}^{+}(x)\phi^{+}_{\ell 0}(x)=0$, and it does not connect to any state of $\mathcal{H}_{\ell}^{-}$, this is why we have excluded $n=0$ in Eq. (31).
In the case that SUSY is broken, none of the equations $\hat{\mathcal{D}}_{\ell}^{\pm}(x)\phi^{\pm}_{\ell 0}(x)=0$ result in a normalizable state. This excludes the existence of zero-energy bound states, while we have a strict isospectrality between the two partner Hamiltonians $\mathcal{H}_{\ell}^{\pm}$, meaning that $\mathcal{E}^{-}_{\ell n}=\mathcal{E}^{+}_{\ell n}$ for all $n\geq 1$. As a result, in each case, the eigenvalues and eigenfunctions of either of the partner Hamiltonians determine the eigenvalues and eigenfunctions of the other one.
The isospectrality between the spectra of two potentials $U_{\ell}^{+}$ and $U_{\ell}^{-}$ originates from the fact that these potentials are SUSY partners, and the singular behaviours of the potentials (at $r=0$ in the case of BH) do not affect the isospectrality.